# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ Appendix A Sufficient conditions for C1
In this section, we provide sufficient conditions for the invertibility of $\mathcal{T}_{P}$ in Condition C1. The key idea is that $\mathcal{T}_{P}$ differs from the identity operator by a compact operator under mild conditions. As a result, $\mathcal{T}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (Conway, 1994).
@@
[…]
where the kernel is given by $K_{P}(a^{\prime},s^{\prime}\mid a,s):=\frac{\pi(a^{\prime}\mid s^{\prime})}{b_{P}(a^{\prime}\mid s^{\prime})}\,\frac{dP(S_{1}=s^{\prime}\mid A_{0}=a,S_{0}=s)}{dP(S_{0}=s^{\prime})}.$ Condition A1 holds with $H=L^{\infty}(\lambda)$ when the state and action spaces are compact subsets of $\mathbb{R}^{d}$, and the kernel function $K_{P}$ is continuous and bounded. In this case, the kernel is square-integrable with respect to the product measure $P_{A_{0},S_{0}}\otimes P_{A_{0},S_{0}}$, so $\mathcal{K}_{P}$ defines a Hilbert–Schmidt operator on $L^{2}(P_{A_{0},S_{0}})$, and is therefore compact (Conway, 1994). Related conditions for the nonparametric well-posedness of the integral equation for the $Q$-function were proposed in Section 3.2 of Chen and Qi (2022). Notably, Condition A1 becomes even less stringent under stronger semiparametric restrictions imposed through $H$, and holds trivially when $H$ is finite-dimensional.
@@
Assume A1 holds at $P\in\mathcal{P}$. Then, the range $\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$ and there exists an element $q_{P,H}\in H$ satisfying (8). Moreover, if A2 also holds, then $\mathcal{T}_{P}\colon(\overline{H}_{P},\|\cdot\|_{P})\rightarrow L^{2}(P_{A_{0},S_{0}})$ has a bounded inverse on its range, and $q_{P,H}=\mathcal{T}_{P}^{-1}(\mu_{P,H})$ is the unique solution.
Condition A1 ensures that the Bellman integral operator $\mathcal{T}_{P}$ can be continuously extended to a map from $\overline{H}_{P}$ to $\overline{H}_{P}$, and this map differs from the identity by a compact operator. Consequently, $\mathcal{T}_{P}=I-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (Conway, 1994). The closedness of the range $\mathcal{T}_{0}(\overline{H}_{0})$, guaranteed by Theorem 8, ensures the existence of the Bellman projection $q_{P,H}$ in (8). Condition A2 further guarantees uniqueness, and holds whenever $\gamma<\|\mathcal{K}_{P}\|_{P}^{-1}$. By the compactness of $\mathcal{K}_{P}$, the spectrum of $\mathcal{T}_{P}$ is countable with $\gamma^{-1}$ as the only possible accumulation point. Hence, the Bellman projection $q_{P,H}$ is unique for almost all discount factors $\gamma$.
@@
Appendix B Model selection with adaptive debiased machine learning
B.1 General Approach
Selecting an appropriate working model is challenging and can compromise inference due to model misspecification bias. Adaptive debiased machine learning (ADML) (van der Laan et al., 2023) provides a unified framework that combines debiased estimation with data-driven model selection to construct superefficient estimators of smooth functionals, adapting to the structure of the nuisance components. By learning model assumptions or feature representations directly from data, ADML facilitates valid inference while mitigating misspecification bias. The calibrated plug-in estimator introduced in the previous section is a special case of ADML. In this section, we extend the ADML framework to the MDP setting, showing how semiparametric DRL can be combined with model selection to construct estimators that adapt to the functional form of the $Q$-function $q_{0}$, going beyond calibration alone. We refer to this extension as Adaptive DRL (ADRL).
For example, the working model $H_{n}$ could be selected via cross-validated FQI over a sieve of models, i.e., a sequence of increasingly complex classes $H_{1}\subset H_{2}\subset H_{3}\subset\dots\subset H_{\infty}:=H$, where $H$ is a correctly specified model containing $q_{0}$. A plausible oracle submodel $H_{0}$ is the smallest correctly specified class in the sieve that contains $q_{0}$, which can feasibly be approximated via cross-validation. Alternatively, $H_{n}$ could result from a variable selection procedure or a learned feature transformation, with $H_{0}$ corresponding to the asymptotically selected variables or limiting transformation. Data-adaptive methods for learning state-action feature representations have been proposed in Pritz et al. (2021) and Pavse and Hanna (2024). Such transformations can also be derived directly from the fitted FQI model $q_{n,H_{n}}^{\pi}$, for example, by one-hot encoding the leaf nodes of trees in a random forest or gradient-boosted tree model, as in Section 3.1 of He et al. (2014).
Example 4 (Domain adaptation and confounding-robust data fusion).
Continuing from Example 3, suppose $S_{t}=(\widetilde{S}_{t},Z)$, where $Z=1$ denotes a randomized experiment with limited data and $Z=0$ denotes abundant but confounded observational data. To improve efficiency, we augment the unbiased experimental data with biased observational data. In Appendix E, we extend the ADML framework of van der Laan et al. (2024d), originally developed for cross-sectional data fusion, to the MDP setting. Following van der Laan et al. (2024d), we may use the Highly Adaptive Lasso (Benkeser and Van Der Laan, 2016) to learn a model class $\mathcal{H}_{n}$ that constrains the difference between the experimental and observational $Q$-functions, i.e., $q_{0}(a,s,1)-q_{0}(a,s,0)$. As shown in Example 3, these constraints enable confounding-robust information sharing across domains while avoiding misspecification by not imposing fixed structural assumptions. Crucially, without such constraints, no efficiency gain is possible from observational data. Learning them from data is therefore essential for improving efficiency while preserving nonparametric validity. See Appendix E for details.
B.2 Asymptotic theory
Theorem 9 (Second-order model approximation bias).
where $H_{n,0}:=H_{n}\oplus H_{0}$ is the direct sum linear model.
To further clarify these conditions, suppose the working model $H_{n}:=H_{\phi_{n}}$ and the oracle model $H_{0}:=H_{\phi_{0}}$ are induced by feature transformations. For a transformation $\phi:\mathcal{A}\times\mathcal{Z}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}$. The combined model $H_{n,0}$ is given by $H_{(\phi_{n},\phi_{0})}$, where $(\phi_{n},\phi_{0})$ denotes the feature map formed by stacking $\phi_{n}$ and $\phi_{0}$. Theorem 4 implies that the approximation bias vanishes if the nuisance functions derived from $\phi_{n}$ and $(\phi_{n},\phi_{0})$ converge to those derived from $\phi_{0}$. The special case $\phi_{n}=q_{n}$ and $\phi_{0}=q_{0}$ recovers Theorem 5 as a corollary. In Lemma 14 (Appendix H.1), we show that with features $X$ and outcome $Y$, the $L^{2}(P_{0})$ error of estimating $E_{0}[Y\mid\phi_{0}(X)]$ using either $E_{0}[Y\mid\phi_{n}(X),\mathcal{D}_{n}]$ or $E_{0}[Y\mid\phi_{n}(X),\phi_{0}(X),\mathcal{D}_{n}]$ is bounded by the feature approximation error $\sqrt{\int\|\phi_{n}(x)-\phi_{0}(x)\|_{\mathbb{R}^{m}}^{2}P_{0,X}(dx)}$. A sufficient condition for this bound is that the map $(t_{1},t_{2})\mapsto E_{0}[Y\mid\phi_{n}(X)=t_{1},\phi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely Lipschitz continuous, generalizing Condition D2.
$\varphi_{P,H}(s,a,y,s^{\prime}):=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{y+\gamma V^{\pi}(q_{P,H}^{\pi})(s^{\prime})-q_{P,H}^{\pi}(a,s)\right\}+m(s,a,q_{P,H})-\Psi_{H}(P),$
which aligns with the influence function in Theorem 1 for $P\in\mathcal{P}_{H}$. Let $\varphi_{n,H_{n}}$ denote the estimator of the influence function $\varphi_{0,H_{n}}$, obtained by plugging in our nuisance estimators.
  1. (C4)
  2. (C5)
    Nuisance estimation rate: $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\|_{P_{0}}\|\mathcal{T}_{0}(\lambda_{n,H_{n}})-\mathcal{T}_{0}(\lambda_{0,H_{n}})\|_{P_{0}}=o_{p}(n^{-1/2})$.
  3. (C6)
  4. (C7)
Theorem 10.
Together, C4 and C5 imply that $\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})=(P_{n}-P_{0})\varphi_{0,H_{n}}+o_{p}(n^{-\frac{1}{2}})$, so that $\psi_{n,H_{n}}$ is debiased for the working parameter $\Psi_{H_{n}}(P_{0})$. Conditions C6 and C7, which ensures that data-driven model selection preserves the validity of the debiased machine learning estimator, appear in prior works on ADML (van der Laan et al., 2023, 2024d). Condition C6 is an asymptotic stability condition requiring the EIF for the learned model $H_{n}$ to converge to the EIF for the oracle submodel $H_{0}$, which necessitates that $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $q_{0,H_{n}}$ are asymptotically consistent with their oracle counterparts. Condition C7 ensures the parameter approximation bias satisfies $\Psi_{n}(P_{0})-\Psi(P_{0})=o_{p}(n^{-\frac{1}{2}})$ in view of Theorem 4.
@@ Appendix C Additional details for Section 5
C.1 Empirical Bellman calibration of Alg. 2
@@
Suppose that $f_{n}^{*}$ is the fixed point isotonic regression solution to the calibrated fitted Q-iteration algorithm in Alg. 2, such that:
$f_{n}^{*}=\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\{Y_{0,i}+\gamma V_{f_{n}^{*}\circ q_{n}}^{\pi}(S_{1,i}))-f(q_{n}(A_{0,i},S_{0,i}))\}^{2}.$
@@
The proof follows from Lemma 4 in Van Der Laan et al. (2023) with minor notational changes. Recall that $f_{n}^{*}$ is the unique càdlàg piecewise constant solution of the isotonic regression problem with jumps occurring only at observed values of $q_{n}$. For any transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, we claim that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is monotone nondecreasing for $\varepsilon$ sufficiently close to zero. To see this, note that $f_{n}^{*}$ is a step function with only finitely many jumps. As a consequence, $h\circ f_{n}^{*}$ is also a step function with the same jump points as $f_{n}^{*}$. By taking $\varepsilon$ close enough to zero, we can guarantee that the maximum jump size of $\varepsilon(h\circ f_{n}^{*})$ is smaller than the minimum jump size of $f_{n}^{*}$. For all $\varepsilon$ sufficiently close to zero, it must then be the case that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is also monotone nondecreasing and, thus, an element of $\mathcal{F}_{\text{iso}}$. Since $f_{n}^{*}$ is the empirical risk minimizer over $\mathcal{F}_{\text{iso}}$, we must have that
$\frac{d}{d\varepsilon}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V_{f_{n}^{*}\circ q_{n}}^{\pi}(S_{1,i}))-(f_{n}^{*}+\varepsilon(h\circ f_{n}^{*}))(q_{n}(A_{0,i},S_{0,i}))\right\}^{2}\bigg{|}_{\varepsilon=0}=0,$
@@ C.3 Bellman Calibration with Cross-Fitting
Algorithm 3 presents a cross-fitted variant of isotonic Bellman calibration (Algorithm 2). We note that the isotonic calibration step in Algorithm 2 should not itself be cross-fitted. Instead, isotonic regression should be applied using the cross-fitted estimates obtained by pooling the out-of-fold predictions. Importantly, this additional fitting step does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (van der Laan et al., 2024c).
Algorithm 3 Isotonic Bellman Calibration with Cross-Fitting
@@
The procedure begins with an initial collection of Q-function estimators $\{q_{n}^{(-i)}\}_{i=1}^{n}$, each trained on a subsample that excludes the $i$th observation to preserve independence. These out-of-fold predictions serve as the input to an iterative calibration procedure that updates each fold-specific Q-function through composition with an isotonic regression fit. At each iteration, a global isotonic calibrator $f_{n}^{(k+1)}\in\mathcal{F}_{\text{iso}}$ is trained to regress the pseudo-outcomes $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(-i,k)})(S_{1,i})$ on the out-of-fold predictions $q_{n}^{(-i)}(A_{0,i},S_{0,i})$. The updated estimate $q_{n}^{*(-i,k+1)}$ is then defined as the composition $f_{n}^{(k+1)}\circ q_{n}^{(-i)}$. The process continues until convergence in $L_{2}(P_{n})$, at which point the final cross-fitted calibrated estimators $\{q_{n}^{*(-i)}\}_{i=1}^{n}$ are returned. This calibration procedure leverages the entire dataset while preserving fold-level independence in each update, thereby recovering full-sample efficiency without violating the theoretical guarantees of cross-fitted debiased machine learning. We refer the reader to van der Laan et al. (2024a) and van der Laan et al. (2024c) for additional examples of calibration on cross-fitted estimates. Importantly, this additional fitting from calibration does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (van der Laan et al., 2024c; Rabenseifner et al., 2025).
@@ C.4 Discussions of conditions of Theorem 7
Condition D1 ensures pathwise differentiability of $\Psi_{q_{0}}$, $\Psi_{q_{n}^{*}}$, and $\Psi_{(q_{n}^{*},q_{0})}$, and requires overlap only in the lower-dimensional feature space induced by the $Q$-function. It is therefore significantly weaker than the condition required for pathwise differentiability of $\Psi$ in the full nonparametric model. Condition D3 holds if the calibrated estimator $q_{n}^{*}$ is $o_{p}(n^{-1/4})$-consistent for $q_{0}$ in $L^{2}(P_{0})$, and if the best approximation $q_{0,q_{n}^{*}}^{\pi}$ to $q_{0}$ given $q_{n}^{*}$ satisfies $\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}}^{\pi})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}=o_{p}(n^{-1/4})$. Together, these imply both $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(n^{-1/4})$ and the required approximation rate in the transformed space. General results on isotonic calibration suggest that $q_{n}^{*}$ converges at least as fast—and possibly faster—than the initial estimator $q_{n}$, up to an asymptotically negligible error of order $n^{-1/3}$ (van der Laan and Alaa, 2025). Condition D4 is satisfied if (i) $\|\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\|_{P_{0}}=o_{p}(1)$, and (ii) the difference lies in a Donsker class or if the initial estimator $q_{n}$ is estimated via sample splitting or cross-fitting (van der Laan et al., 2024a; Rabenseifner et al., 2025). Under boundedness and D2, the first condition holds whenever $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(1)$. This empirical process condition holds under mild conditions when Alg. 3 is used.

### v4

@@ Appendix A Sufficient conditions for C1
In this section, we provide sufficient conditions for the invertibility of $\mathcal{T}_{P}$ in Condition C1. The key idea is that $\mathcal{T}_{P}$ differs from the identity operator by a compact operator under mild conditions. As a result, $\mathcal{T}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (conway1994course).
@@
$\mathcal{K}_{P}(h)(a,s)=\int h(a^{\prime},s^{\prime})\,K_{P}(a^{\prime},s^{\prime}\mid a,s)\,dP_{A_{0},S_{0}}(a^{\prime},s^{\prime}),$
where the kernel is given by $K_{P}(a^{\prime},s^{\prime}\mid a,s):=\frac{\pi(a^{\prime}\mid s^{\prime})}{b_{P}(a^{\prime}\mid s^{\prime})}\,\frac{dP(S_{1}=s^{\prime}\mid A_{0}=a,S_{0}=s)}{dP(S_{0}=s^{\prime})}.$ Condition A1 holds with $H=L^{\infty}(\lambda)$ when the state and action spaces are compact subsets of $\mathbb{R}^{d}$, and the kernel function $K_{P}$ is continuous and bounded. In this case, the kernel is square-integrable with respect to the product measure $P_{A_{0},S_{0}}\otimes P_{A_{0},S_{0}}$, so $\mathcal{K}_{P}$ defines a Hilbert–Schmidt operator on $L^{2}(P_{A_{0},S_{0}})$, and is therefore compact (conway1994course). Related conditions for the nonparametric well-posedness of the integral equation for the $Q$-function were proposed in Section 3.2 of chen2022well. Notably, Condition A1 becomes even less stringent under stronger semiparametric restrictions imposed through $H$, and holds trivially when $H$ is finite-dimensional.
@@
Assume A1 holds at $P\in\mathcal{P}$. Then, the range $\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$ and there exists an element $q_{P,H}\in H$ satisfying (7). Moreover, if A2 also holds, then $\mathcal{T}_{P}\colon(\overline{H}_{P},\|\cdot\|_{P})\rightarrow L^{2}(P_{A_{0},S_{0}})$ has a bounded inverse on its range, and $q_{P,H}=\mathcal{T}_{P}^{-1}(r_{P,H})$ is the unique solution.
Condition A1 ensures that the Bellman integral operator $\mathcal{T}_{P}$ can be continuously extended to a map from $\overline{H}_{P}$ to $\overline{H}_{P}$, and this map differs from the identity by a compact operator. Consequently, $\mathcal{T}_{P}=I-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (conway1994course). The closedness of the range $\mathcal{T}_{0}(\overline{H}_{0})$, guaranteed by Theorem 7, ensures the existence of the Bellman projection $q_{P,H}$ in (7). Condition A2 further guarantees uniqueness, and holds whenever $\gamma<\|\mathcal{K}_{P}\|_{P}^{-1}$. By the compactness of $\mathcal{K}_{P}$, the spectrum of $\mathcal{T}_{P}$ is countable with $\gamma^{-1}$ as the only possible accumulation point. Hence, the Bellman projection $q_{P,H}$ is unique for almost all discount factors $\gamma$.
@@
Appendix B Additional details on nuisance estimation
Algorithm 2 Fitted Q-Iteration
0:  Data, Function class $H$, number of iterations $K$;
1:  Initialize $q_{n,H}^{\pi,(0)}:=0$;
2:  for $k=0,1,\ldots,K-1$ do
3:   Set value function $V_{n,H}^{\pi,(k)}:s\mapsto\int q_{n,H}^{\pi,(k)}(a^{\prime},s)\pi(a^{\prime}\mid s)da^{\prime};$
4:   Update $q_{n,H}^{\pi,(k+1)}\in H$ by estimating:
5:  end for
6:  Set $q_{n,H}:=q_{n,H}^{\pi,(K)}$;
7:  return $q_{n,H}$;
@@ Appendix C Additional details for Section 4
C.1 Empirical Bellman calibration of Alg. 1
@@
Suppose that $f_{n}^{*}$ is the fixed point isotonic regression solution to the calibrated fitted Q-iteration algorithm in Alg. 1, such that:
$f_{n}^{*}=\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\{Y_{0,i}+\gamma V^{\pi}(f_{n}^{*}\circ q_{n})(S_{1,i}))-f(q_{n}(A_{0,i},S_{0,i}))\}^{2}.$
@@
The proof follows from Lemma 4 in van2023causal with minor notational changes. Recall that $f_{n}^{*}$ is the unique càdlàg piecewise constant solution of the isotonic regression problem with jumps occurring only at observed values of $q_{n}$. For any transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, we claim that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is monotone nondecreasing for $\varepsilon$ sufficiently close to zero. To see this, note that $f_{n}^{*}$ is a step function with only finitely many jumps. As a consequence, $h\circ f_{n}^{*}$ is also a step function with the same jump points as $f_{n}^{*}$. By taking $\varepsilon$ close enough to zero, we can guarantee that the maximum jump size of $\varepsilon(h\circ f_{n}^{*})$ is smaller than the minimum jump size of $f_{n}^{*}$. For all $\varepsilon$ sufficiently close to zero, it must then be the case that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is also monotone nondecreasing and, thus, an element of $\mathcal{F}_{\text{iso}}$. Since $f_{n}^{*}$ is the empirical risk minimizer over $\mathcal{F}_{\text{iso}}$, we must have that
$\frac{d}{d\varepsilon}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V^{\pi}(f_{n}^{*}\circ q_{n})(S_{1,i}))-(f_{n}^{*}+\varepsilon(h\circ f_{n}^{*}))(q_{n}(A_{0,i},S_{0,i}))\right\}^{2}\bigg|_{\varepsilon=0}=0,$
@@ C.3 Bellman Calibration with Cross-Fitting
Algorithm 3 presents a cross-fitted variant of fitted $Q$-calibration (Algorithm 1). We note that the isotonic calibration step in Algorithm 1 should not itself be cross-fitted. Instead, isotonic regression should be applied using the cross-fitted estimates obtained by pooling the out-of-fold predictions. Importantly, this additional fitting step does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (vanderlaan2024automatic).
Algorithm 3 Isotonic Fitted $Q$-calibration with Cross-Fitting
@@
The procedure begins with an initial collection of Q-function estimators $\{q_{n}^{(-i)}\}_{i=1}^{n}$, each trained on a subsample that excludes the $i$th observation to preserve independence. These out-of-fold predictions serve as the input to an iterative calibration procedure that updates each fold-specific Q-function through composition with an isotonic regression fit. At each iteration, a global isotonic calibrator $f_{n}^{(k+1)}\in\mathcal{F}_{\text{iso}}$ is trained to regress the pseudo-outcomes $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(-i,k)})(S_{1,i})$ on the out-of-fold predictions $q_{n}^{(-i)}(A_{0,i},S_{0,i})$. The updated estimate $q_{n}^{*(-i,k+1)}$ is then defined as the composition $f_{n}^{(k+1)}\circ q_{n}^{(-i)}$. The process continues until convergence in $L_{2}(P_{n})$, at which point the final cross-fitted calibrated estimators $\{q_{n}^{*(-i)}\}_{i=1}^{n}$ are returned. This calibration procedure leverages the entire dataset while preserving fold-level independence in each update, thereby recovering full-sample efficiency without violating the theoretical guarantees of cross-fitted debiased machine learning. We refer the reader to van2024stabilized and vanderlaan2024automatic for additional examples of calibration on cross-fitted estimates. Importantly, this additional fitting from calibration does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (vanderlaan2024automatic; rabenseifner2025calibration).
@@ C.4 Discussions of conditions of Theorem 6
Condition D1 ensures pathwise differentiability of $\Psi_{q_{0}}$, $\Psi_{q_{n}^{*}}$, and $\Psi_{(q_{n}^{*},q_{0})}$, and requires overlap only in the lower-dimensional feature space induced by the $Q$-function. It is therefore significantly weaker than the condition required for pathwise differentiability of $\Psi$ in the full nonparametric model. Condition D3 holds if the calibrated estimator $q_{n}^{*}$ is $o_{p}(n^{-1/4})$-consistent for $q_{0}$ in $L^{2}(P_{0})$, and if the best approximation $q_{0,q_{n}^{*}}^{\pi}$ to $q_{0}$ given $q_{n}^{*}$ satisfies $\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}}^{\pi})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}=o_{p}(n^{-1/4})$. Together, these imply both $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(n^{-1/4})$ and the required approximation rate in the transformed space. General results on isotonic calibration suggest that $q_{n}^{*}$ converges at least as fast—and possibly faster—than the initial estimator $q_{n}$, up to an asymptotically negligible error of order $n^{-1/3}$ (van2025generalized). Condition D4 is satisfied if (i) $\|\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\|_{P_{0}}=o_{p}(1)$, and (ii) the difference lies in a Donsker class or if the initial estimator $q_{n}$ is estimated via sample splitting or cross-fitting (van2024stabilized; rabenseifner2025calibration). Under boundedness and D2, the first condition holds whenever $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(1)$. This empirical process condition holds under mild conditions when Alg. 3 is used.